Suppose that the production function is given by = 2x0.5. The price of is $3 and the price of is $4. Derive the corresponding VMP and AVP functions. What is MFC? Solve for the profit-maximizing level for input use x.


1
Expert's answer
2021-01-20T14:59:56-0500

Let "P_x=\$3, P_y=\$4."

Then, the total value of the product (TVP) can be written as follows:


"TVP=P_y\cdot y."


The VMP can be found as follows:


"VMP=\dfrac{dTVP}{dx}=P_y\dfrac{d}{dx}(2x^{0.5})=P_y\cdot\dfrac{1}{\sqrt{x}}."


The total factor cost (TFC) can be written as follows:


"TFC=P_x\cdot x."

The profit function can be written as:


"\Pi=TVP-TFC=P_y\cdot y-P_x\cdot x."

By the definition, the marginal factor cost (MFC) equals:


"MFC=\dfrac{dTFC}{dx}=\dfrac{d}{dx}(P_x\cdot x)=P_x=\$3."

Average value product (AVP) can be written as:


"AVP=P_y\cdot2x^{0.5}\cdot\dfrac{1}{x}."

The profit-maximizing level:


"\dfrac{d}{dx}\Pi=0,""P_y\dfrac{dy}{dx}-P_x=0,""P_y\dfrac{d}{dx}(2x^{0.5})-P_x=0,""P_y\cdot\dfrac{1}{\sqrt{x}}-P_x=0,""x=(\dfrac{P_y}{P_x})^2=(\dfrac{\$4}{\$3})^2=1.77"